Calculate the mass of ammonia required to produce 3.01x10^23 Molecules of nitrogen according to the following equation 4NH3+3O2➡️2N2+6H2O


1
Expert's answer
2021-02-22T06:08:44-0500

n(N2) = 3.01x10^23/6.02x10^23 = 0.5 mol

n(NH3) = 0.5*2 = 1 mol

m(NH3) = 1mol*17g/mol = 17 g